OpenStudy (anonymous):

A sequence is defined recursively using the formula f(n + 1) = f(n) – 5 . Which sequence could be generated using this formula? 1, –5, 25, –125, ... 2, 10, 50, 250, ... 3, –2, –7, –12, ... 4, 9, 14, 19, ...

OpenStudy (anonymous):

@A_clan do you know how to do these?

OpenStudy (a_clan):

When you substitute n=1,2,3,.. in the formula, we get f(2) = f(1) -5 f(3)=f(2) -5 f(4)=f(3) -5 and so on which shows that any term in the sequence is 5 less than its predecessor. So for the answer, you need to look for a sequence which is decreasing by 5.